OpenStudy (lgbasallote):

LGBADERIVATIVES: \[\huge G(x) = \frac{1-\cosh x}{1+\cosh x}\]

OpenStudy (ash2326):

Use the division rule, it's pretty easy

OpenStudy (lgbasallote):

quotient rule immediately @ash2326 ? cant it be easier???

OpenStudy (lgbasallote):

it looks intimidating

OpenStudy (lgbasallote):

\[\huge \frac{\sinh x(1+\cosh x) - \sinh x(1-\cosh x)}{(1+\cosh x)^2}?\]

OpenStudy (ash2326):

It's easy since \[\frac{d}{dx} (\cos hx)=\sinh x \] \[\frac{(1+\cosh x)\times (-\sinh x)-(1-\cosh x)\times (\sinh x)}{(1+\cosh x)^2} \] just simplification now

OpenStudy (lgbasallote):

\[\huge \frac{\sinh x(2\cosh x)}{(1+\cosh x)^2}?\]

OpenStudy (ash2326):

just \[\frac{-2\sinh x}{(1+\cosh x)^2}\]

OpenStudy (lgbasallote):

why negative?

OpenStudy (lgbasallote):

oh wait yes

OpenStudy (ash2326):

ha ha

OpenStudy (lgbasallote):

i forgot it was 1-coshx